solve: 6r^3 + 7r^2 = 3r (please show me step by step, it would be very helpful)
Answer by lwsshak3(11628)

solve: 6r^3 + 7r^2 = 3r (please show me step by step, it would be very helpful)
**
6r^3 + 7r^2 = 3r
set equation=0
6r^3 + 7r^2 - 3r=0
factor out r
r(6r^2+7r-3)=0
factor trinomial
r(2r+3)(3r-1)=0
solve for zeros or roots
r=0
..
2r+3=0
r=-3/2
..
3r-1=0
r=1/3